6. Using the same sample data as in question 3 above (ie on number of plots sold overa period of six months), Compute the variance of the sample using the formula a. (n -D bRepeat part a), but this time use the alternative formula for calculating sanmple variance (n-1) c. Find the standard deviation of the sample dDescribe a situation where the Coefficient of variation would be a more useful istical measure than Standard deviation. the coofficient of variation for the sample 7. You are on an appointment panel and have the following information on the exam renuli of two potential candidates Candidate A Scored 53 in a test for which μ-SO and σ-10 Candidate B: Scored 53 in a test for which μ 50 and σ-5 Assuming there is no other difference between the rwo applicants, which candidate would yo fovour prior to the interview itself [Hint compare their -scores)
Answer #1

7) Candidate A

Given Score = 53,

mu = 50

sigma = 10

Z-Score = rac{x-mu}{}sigma = 53-50 10 = 0.3

for candidate B

given score = 53

mu = 50

sigma = 5

Z-Score = rac{x-mu}{}sigma = 53-50 = 0.6

we select candidate B since he got more Z-score compared to candidate A.